[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Aug 14, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Aug 14 09:20:59 UTC 2017


HSU Shift Log for Mon Aug 14 2017 05:20:52 GMT-0400 (EDT)

Shift: A

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: Microwave imagery through 0230Z are available in ATCF for  
Gert. However, given the plethora of MW satellites orbiting Earth, there  
seems to be a unusual dearth of any MI imagery on either the FNMOC and TC  
web pages, and in the ATCF for Gert as of 0900Z. MI for Invest 91L seems to  
be okay on both web pages.

N-AWIPS issues: Atlantic N-AWIPS (bottom/left of ATCF) screen went  
completely 'green' around 0330Z; at 0400Z, the main Atlantic N-AWIPS screen  
(to right of ATCF) went completely 'pink'. After multiple attempts, was  
finally able to reboot both work stations, which corrected the problems.

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170814/e757671b/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list